OpenStudy (anonymous):

Solve the following quadratic equation by using the square root property (x - 2)2 = 5 Can somebody review me on how to do this again?

OpenStudy (anonymous):

is that the right question?

OpenStudy (anonymous):

(x - 2)2 = 5 taking squareroot of both sides x - 2 = ± √5 x = 2 ± √5 so x = 2 + √5 and x = 2 - √5